# Bangkok 10330, Thailand

### v1

@@ 1 Introduction
The false vacuum decay is one of the important phenomena in quantum theories, already at the level of quantum mechanics, with interesting applications to a variety of areas, such as phase transitions, non perturbative dynamics of quantum field theories and cosmology. A standard approach to study the decay is by saturating the Euclidean path integral with a classical configuration of finite action that extrapolates between the false and true vacuum corresponding to a bubble of the true vacuum inside the false vacuum, the so-called Bounce Coleman:1977py ; Coleman:1980aw ; Hawking:1981fz . The study of the Bounce in a scalar quantum field theory was initiated long ago in the thin-wall approximation and the decay width was computed by extremizing the action with respect to the position of the wall Coleman:1977py . The analysis was also generalised in the presence of gravity and explicit expressions were given for the transitions from a de Sitter (dS) to Minkowski and from a Minkowski to anti-de Sitter (AdS) vacuum Coleman:1980aw . It was later realised that the thin-wall approximation can be interpreted as a 3-dimensional Euclidean world-volume of a brane separating the false and the true vacuum. The extremization condition can be replaced by imposing the Israel matching conditions on the brane Israel:1966rt , leading to the same result Garfinkle:1989mv ; Chamblin:1999ya (for a recent review see e.g. Gregory:2023eos and references therein).
@@
In this work, we study the false vacuum decay in the presence of compact dimensions, restricting to the simplest illustrative case of a single one. The main complication comes from the symmetry of the Bounce which cannot be spherical $O(d)$ but $O(d-1)\times U(1)$, describing a bubble geometry of a $S^{d-2}$ sphere times a circle $S^{1}$ with their respective radii depending on the radial coordinate. This is parameterised by two functions of the radial coordinate, associated to the radius of $S^{d-2}$ and to the radius of the compact dimension. It turns out that the extremization procedure of the Bounce as in CdL becomes ambiguous while the Israel matching conditions are well defined and lead to simple explicit expressions for the Bounce and for the transition probabilities in the cases of dS to Minkowski and Minkowski to AdS vacua.
@@ 2 Vacuum Decay in Theories with $O(d)$ Symmetry
where $\mathrm{d}\Omega_{d-1}^{2}=g_{ij}\mathrm{d}y^{i}\mathrm{d}y^{j}$ represents the surface element of a $(d-1)$-dimensional unit sphere of radius $H^{-1}$. In the following we shall denote derivatives with respect to $\xi$ as
@@ 2.1 The approach of Coleman and de Luccia (CdL)
  • Mink $\rightarrow$ AdS.   Setting $V(\phi_{+})=0$ and $V(\phi_{-})=-\epsilon$ simplifies (36) as
@@ 3 Vacuum Decay in Theories with $O(d-1)\times U(1)$ Symmetry
$\frac{6\dot{\rho}^{2}}{\rho^{2}}+\frac{3\dot{\rho}\dot{\sigma}}{\rho{\sigma}}-\frac{6H^{2}}{\rho^{2}}=\kappa_{d}\bigg{(}\frac{\dot{\phi}^{2}}{2}-V\bigg{)}\,,$ (70)
$\frac{3\ddot{\rho}}{\rho}+\frac{\ddot{\sigma}}{{\sigma}}+\frac{4\dot{\rho}\dot{\sigma}}{\rho{\sigma}}+\frac{3\dot{\rho}^{2}}{\rho^{2}}-\frac{2H^{2}}{\rho^{2}}=-\kappa_{d}\bigg{(}\frac{\dot{\phi}^{2}}{2}+V\bigg{)}\,,$ (71)
$\frac{3\ddot{\rho}}{\rho}+\frac{3\dot{\rho}^{2}}{\rho^{2}}-\frac{6H^{2}}{\rho^{2}}=-\kappa_{d}\bigg{(}\frac{\dot{\phi}^{2}}{2}+V\bigg{)}\,,$ (72)
@@ 3.1 Instanton geometry
3.2 Bounce and complications of the CdL approach.
@@ 4 Two explicit examples of vacuum transitions
Another important property of the above bounce action is that it only depends on $H$ and $\gamma_{-}$ via the combination $R_{0}\equiv\sqrt{2}\gamma_{-}H$ which is the radius of the compact dimension in the true vacuum (inside the bubble), see (112). Its value at the origin $R_{0}\equiv(\rho\sigma)|_{\xi=0}$ also coincides with the radius at $\xi=\bar{\xi}$, namely $R_{\bar{\xi}}\equiv(\rho\sigma)|_{\xi=\bar{\xi}}=R_{0}$. Note that in the gravity decoupled limit, $B$ is proportional to the only dimensionless quantity of $S_{1},\epsilon,R_{0}$ which is linear in $R_{0}$ and analytic in $S_{1}$, which appears in the lowest integer power. This is to be contrasted with the dimensionless ratio $S_{1}^{5}/\epsilon^{4}$ obtained in the non compact case (41). This is obtained by replacing $R_{0}$ by $S_{1}\over\epsilon$ which has dimension of length, up to a numerical constant. On the other hand, in the limit of vanishing radius, one recovers the 4-dimensional quantity $S_{1}^{4}/\epsilon^{3}$ upon rescaling $S_{1}$ and $\epsilon$ by $1/R_{0}$ as dictated by standard dimensional reduction.
@@ 4.2 Minkowski $\rightarrow$ AdS
which leads to a a bounce action with the following expansion in the gravity decoupling limit
@@
We finally notice that, when introducing the dimensionless variable $\hat{\kappa}_{5}$, the allowed range (149) of $S_{1}$ translates into $0\leq\hat{\kappa}_{5}\leq\frac{3}{2}$, which does not restrict the domain of interest $0\leq\hat{\kappa}_{5}\leq 1$. Hence, we can once again compare the plots for $\mathcal{B}(\hat{\kappa}_{d})$, in the range $\hat{\kappa}_{d}\in[0,1]$, for the Minkowski to AdS transition discussed in subsection 2.1 for $d=4,5$ and above for $d=5$ with a compact dimension.
@@ 5 Conclusions
The interesting question of going beyond the thin wall approximation remains open, in relation to the question of when it breaks down and when the Hawking-Moss classical transition starts dominating.
@@ Acknowledgements.
IA is supported by the Second Century Fund (C2F), Chulalongkorn University. The work of DB and HI has been supported by Thailand NSRF via PMU-B, grant number B13F670063.

### v2

@@ 1 Introduction
The false vacuum decay is one of the important phenomena in quantum theories, already at the level of quantum mechanics, with interesting applications to a variety of areas, such as phase transitions, non perturbative dynamics of quantum field theories and cosmology. A standard approach to study the decay is by saturating the Euclidean path integral with a classical configuration of finite action that extrapolates between the false and true vacuum corresponding to a bubble of the true vacuum inside the false vacuum, the so-called Bounce Coleman:1977py ; Coleman:1980aw ; Hawking:1981fz . The study of the Bounce in a scalar quantum field theory was initiated long ago in the thin-wall approximation and the decay width was computed by extremising the action with respect to the position of the wall Coleman:1977py . The analysis was also generalised in the presence of gravity and explicit expressions were given for the transitions from a de Sitter (dS) to Minkowski and from a Minkowski to anti-de Sitter (AdS) vacuum Coleman:1980aw . It was later realised that the thin-wall approximation can be interpreted as a 3-dimensional Euclidean world-volume of a brane separating the false and the true vacuum. The extremisation condition can be replaced by imposing the Israel matching conditions on the brane Israel:1966rt , leading to the same result Garfinkle:1989mv ; Chamblin:1999ya (for a recent review see e.g. Gregory:2023eos and references therein).
@@
In this work, we study the false vacuum decay in the presence of compact dimensions, restricting to the simplest illustrative case of a single one [^1] For interesting related work involving the presence of non-dynamical compact directions see for example Kleban:2011cs ; Apruzzi:2019ecr ; Apruzzi:2021nle . For non-Euclidean approaches to the vacuum decay problem see for example Garcia-Compean:2021syl . . The main complication comes from the symmetry of the Bounce which cannot be spherical $O(d)$ but $O(d-1)\times U(1)$, describing a bubble geometry of a $S^{d-2}$ sphere times a circle $S^{1}$ with their respective radii depending on the radial coordinate. This is parameterised by two functions of the radial coordinate, associated to the radius of $S^{d-2}$ and to the radius of the compact dimension. It turns out that the extremisation procedure of the Bounce as in CdL becomes ambiguous while the Israel matching conditions are well defined and lead to simple explicit expressions for the Bounce and for the transition probabilities in the cases of dS to Minkowski and Minkowski to AdS vacua.
@@ 2 Vacuum Decay in Theories with $O(d)$ Symmetry
where $\mathrm{d}\Omega_{d-1}^{2}=g_{ij}\mathrm{d}y^{i}\mathrm{d}y^{j}$ represents the surface element of a $(d-1)$-dimensional sphere of radius $H^{-1}$. In the following we shall denote derivatives with respect to $\xi$ as
@@ 2.1 The approach of Coleman and de Luccia (CdL)
  • Minkowski $\rightarrow$ AdS.   Setting $V(\phi_{+})=0$ and $V(\phi_{-})=-\epsilon$ simplifies (36) as
@@ 3 Vacuum Decay in Theories with $O(d-1)\times U(1)$ Symmetry
$\frac{6\dot{\rho}^{2}}{\rho^{2}}+\frac{3\dot{\rho}\dot{\sigma}}{\rho{\sigma}}-\frac{3}{\alpha^{2}\rho^{2}}=\kappa_{d}\bigg{(}\frac{\dot{\phi}^{2}}{2}-V\bigg{)}\,,$ (70)
$\frac{3\ddot{\rho}}{\rho}+\frac{\ddot{\sigma}}{{\sigma}}+\frac{4\dot{\rho}\dot{\sigma}}{\rho{\sigma}}+\frac{3\dot{\rho}^{2}}{\rho^{2}}-\frac{1}{\alpha^{2}\rho^{2}}=-\kappa_{d}\bigg{(}\frac{\dot{\phi}^{2}}{2}+V\bigg{)}\,,$ (71)
$\frac{3\ddot{\rho}}{\rho}+\frac{3\dot{\rho}^{2}}{\rho^{2}}-\frac{3}{\alpha^{2}\rho^{2}}=-\kappa_{d}\bigg{(}\frac{\dot{\phi}^{2}}{2}+V\bigg{)}\,,$ (72)
@@ 3.1 Instanton geometry
3.2 Bounce and complications of the CdL approach
@@ 4 Two explicit examples of vacuum transitions
Another important property of the above bounce action is that it only depends on $H$ and $\gamma_{-}$ via the combination $R_{0}\equiv\sqrt{2}\gamma_{-}H$ which is the radius of the compact dimension in the true vacuum (inside the bubble), see (112). Its value at the origin $R_{0}\equiv(\rho\sigma)|_{\xi=0}$ also coincides with the radius at $\xi=\bar{\xi}$, namely $R_{\bar{\xi}}\equiv(\rho\sigma)|_{\xi=\bar{\xi}}=R_{0}$. Note that in the gravity decoupled limit, $B$ is proportional to the only dimensionless quantity of $S_{1},\epsilon,R_{0}$ which is linear in $R_{0}$ and analytic in $S_{1}$, appearing in the lowest integer power. This is to be contrasted with the dimensionless ratio $S_{1}^{5}/\epsilon^{4}$ obtained in the non compact case (41). This is obtained by replacing $R_{0}$ by $S_{1}/\epsilon$ which has dimension of length, up to a numerical constant. On the other hand, in the limit of vanishing radius, one recovers the 4-dimensional quantity $S_{1}^{4}/\epsilon^{3}$ upon rescaling $S_{1}$ and $\epsilon$ by $1/R_{0}$ as dictated by standard dimensional reduction.
@@
[…]
which is incompatible with constant $\sigma(\xi)$ for positive $V$ (the case of negative $V$ corresponding to a transition between two AdS vacua is more involved). Notice however that in the Lorentzian case the two terms on the left-hand side above come with a relative minus sign and there is no obstruction; one obtains a 5d de Sitter solution where the compact and non-compact dimensions expand exponentially with time Anchordoqui:2023etp .
In turn, the obstruction arising for solutions with constant $\sigma(\xi)$ can also be regarded as the impossibility of changing the topological structure of the bounce from $O(4)\times U(1)$ to $O(5)$ by simply taking the limit of the $U(1)$-radius to infinity, as one may naively expect. We notice that this limit exists in the case of flat spatial sections, i.e. $S^{3}$ is replaced by $\mathbb{R}^{3}$, leading to a locally dS5 or AdS5 solution (depending on the sign of the potential $V$) with one compact space dimension. In this case, there is no topological obstruction but the solution has not finite action.
Figure 1: Behaviour of the dimensionless function $\mathcal{B}(\hat{\kappa}_{d})$ describing the gravitational corrections in the dS to Minkowski transition, normalised to the result in the absence of gravity, as a function of the dimensionless parameter $\hat{\kappa}_{d}$ defined in (37). The plot compares the results of the $d=4,5$ non-compact case obtained in section 2 with those of $d=5$ with one compact dimension studied here.
@@ 4.2 Minkowski $\rightarrow$ AdS
which leads to a bounce action with the following expansion in the gravity decoupling limit
@@
We finally notice that, when introducing the dimensionless variable $\hat{\kappa}_{5}$, the allowed range (150) of $S_{1}$ translates into $0\leq\hat{\kappa}_{5}\leq\frac{3}{2}$, which does not restrict the domain of interest $0\leq\hat{\kappa}_{5}\leq 1$. Hence, we can once again compare, see figure 2, the plots for $\mathcal{B}(\hat{\kappa}_{d})$, in the range $\hat{\kappa}_{d}\in[0,1]$, for the Minkowski to AdS transition discussed in subsection 2.1 for $d=4,5$ and above for $d=5$ with a compact dimension. Note that, as already observed in the dS to Minkowski transition, the curve for the $d=5$ case with one compact dimension does not lie between the $d=4$ and $d=5$ non-compact curves, and does not interpolate between the limits $R_{0}\rightarrow 0$ and $R_{0}\rightarrow\infty$; see discussion above figure 1 and conclusions.
Figure 2: Behaviour of the dimensionless function $\mathcal{B}(\hat{\kappa}_{d})$ describing the gravitational corrections in the Minkowski to AdS transition. The plot compares the results of the $d=4,5$ non-compact case obtained in section 2 with those of $d=5$ with one compact dimension studied here.
@@ 5 Conclusions
[…]
Another interesting question of going beyond the thin wall approximation remains open, in relation to the question of when it breaks down and when the Hawking-Moss classical transition starts dominating.
@@ Acknowledgements.
IA is supported by the Second Century Fund (C2F), Chulalongkorn University. The work of DB and HI has been supported by Thailand NSRF via PMU-B, grant number B13F670063. IA would like to thank Matt Kleban for enlightening correspondence.